# Purified pseudomode model for nonlinear system-bath interactions

### v1

@@ III.2 Explicit pseudomode model for the nonlinear system-bath interaction (2)
As in the analysis of Ref. [79], we assume a spectral decomposition of its positive contribution $C(t\geq 0)$ in the form of
@@
with $w_{l}\in\mathbb{C}$, $\nu_{l}\in\mathbb{R}$, $\gamma_{l}\in\mathbb{R}_{+}$, and the condition $\sum_{l=1}^{N}w_{l}\cong\sigma^{2}$ is satisfied for consistency. Such an exponential decomposition can be realized using several existing algorithms, e.g., the AAA algorithm and the Prony decomposition, see Ref. [82] and the references therein for a comprehensive review. Because of the time-reversal property $C(t)=C^{*}(-t)$, the decomposition in Eq. (15) also gives the decomposition of the negative contribution $C(t\leq 0)$, written as
@@
To match the pseudomode correlation ansatz (13), we combine the two decompositions in Eqs. (15) and (16) to obtain the following approximation
@@
$C_{\text{pm},\gtrless}^{a}(t)=\sum_{l=1}^{N}(\text{Re}\,w_{l}\pm i\text{Im}\,w_{l})e^{-i(\nu_{l}\pm ia)t-(\gamma_{l}+a)\lvert t\rvert},\$ (18)
where $a\in[0,+\infty)$ is a free parameter. The use of the symbol $\cong$ in Eq. (17) highlights that the approximating error is dominated by the one made using the approximation ansatz in Eq. (15) and it requires further justification, which we now provide. Comparing Eq. (18) with Eq. (15) and Eq. (16), we find that
$\begin{array}[]{lll}C_{\text{pm},\gtrless}^{a}(t\gtrless 0)&\cong&C(t\gtrless 0),\\ C_{\text{pm},\gtrless}^{a}(t\lessgtr 0)&\sim&\mathcal{O}(e^{-a\lvert t\rvert})\;,\end{array}$ (19)
defining the intended forward- and backward-time nature of, respectively, $C^{a}_{\text{pm},>}(t)$ and $C^{a}_{\text{pm},<}(t)$, in the limit $a\to+\infty$. In other words, for $t\neq 0$, the overall approximation error $\lvert C(t)-C_{\text{pm}}^{a}(t)|$ arising from the introducing of the free parameter $a$ is of order $\mathcal{O}(e^{-a\lvert t\rvert})$, which vanishes in the limit $a\to+\infty$, justifying $C(t)\cong\lim_{a\to+\infty}C_{\text{pm}}^{a}(t)$ for $t\neq 0$. Further analysis for the case $t=0$ is necessary due to the explicit dependence of the dynamics on $C(0)$ in the nonlinear case analyzed here, as shown at the end of section Sec. III.1. Specifically, we introduced the additional term $C_{\text{zf}}(t)$ in Eq. (17) to correctly reproduce the second moment of $X(0)$, as can be explicitly checked as
@@
In this way, we have thus defined a pseudomode model consisting of $12N+$ 1 zero-temperature pseudomodes, so that the sum of their free correlations $C_{\text{pm}}^{a}(t)$ approximates $C(t)$ with the error dominating by the spectral decomposition in Eq. (15) in the limit $a\to+\infty$. This completes the construction of an explicit pseudomode model for general nonlinear system-bath interactions (2).
III.3 Purified pseudomode model for the special case $Q=X^{n}$
In our previous work [79], we have shown that when the frequency $\Omega$ and decay rate $\Gamma$ of a zero-temperature pseudomode is analytically continued along either of the paths $\Omega\to\Omega\pm ia$ and $\Gamma\to\Gamma+a$, the state of the pseudomode becomes effectively pure when the limit $a\to+\infty$ is reached, leading to an optimization of the pseudomode Hilbert space which can be exploited in numerical simulations. This important result allows us to introduce the concept of purified pseudomodes, which, formally, can be understood as pseudomodes characterizied by the frequency $\Omega\pm i\infty$ and the decay rate $\Gamma+\infty$. Interestingly, the purified pseudomode correlations, obtained by taking the limit $C_{\text{ppm}}^{\pm}(t)=\lim_{a\to+\infty}\lambda^{2}\exp(-i(\Omega\pm ia)t-(\Gamma+a)\lvert t\rvert)=\lambda^{2}\exp(-i\Omega t\mp\Gamma t)\Theta(\pm t)$, correspond to either the positive-time ($+$) or negative-time (--) contributions of the pseudomode correlation, see Eq. (13).
@@ III.3 Purified pseudomode model for the special case $Q=X^{n}$
In Appendix. III.3, we generalize the purification technique in the previous work [79] to the special case $Q=X^{n}$, and show that in the limit $a\to+\infty$ the equation of motion in Eq. (10) becomes the following Lindblad-like master equation
$i\frac{d\rho_{\text{ppm}}(t)}{dt}=[H_{S},\rho_{\text{ppm}}]+\mathcal{L}_{0}\rho_{\text{ppm}}+\mathcal{L}_{\text{int},n}\rho_{\text{ppm}},$ (23)
with the generators $\mathcal{L}_{0}$ and $\mathcal{L}_{\text{int},n}$ defined as
@@
$\begin{array}[]{lll}&\displaystyle\mathcal{L}_{0}[\boldsymbol{\cdot}]=\sum_{l=1}^{N}\Big{[}(\nu_{l}-i\gamma_{l})d_{l+}^{\dagger}d_{l+}[\boldsymbol{\cdot}]-(\nu_{l}+i\gamma_{l})[\boldsymbol{\cdot}]d_{l-}^{\dagger}d_{l-}\Big{]},\\ &\displaystyle\mathcal{L}_{\text{int},n}[\boldsymbol{\cdot}]=X_{+}^{n}s[\boldsymbol{\cdot}]-[\boldsymbol{\cdot}]sX_{-}^{n}+s[\boldsymbol{\cdot}]X_{-,d^{\dagger}}^{n}-X_{+,d}^{n}[\boldsymbol{\cdot}]s\\ &\displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{r=1}^{n-1}\binom{n}{r}\bigg{[}X_{+}^{n-r}s[\boldsymbol{\cdot}]X_{-,d^{\dagger}}^{r}-X_{+,d}^{r}[\boldsymbol{\cdot}]sX_{-}^{n-r}\bigg{]},\end{array}$ (24)
@@
where $\rho_{\text{ppm}}(t)$ denotes the density matrix describing the state of the system $S$ and all purified pseudomodes, and we have defined the operators $X_{\pm}=X_{\pm,d}+X_{\pm,d^{\dagger}}$ with $X_{\pm,d}=\sum_{l=1}^{N}\lambda_{l\pm}d_{l\pm}$ and $X_{\pm,d^{\dagger}}=\sum_{l=1}^{N}\lambda_{l\pm}d_{l\pm}^{\dagger}$ their components to simplify the notation. The initial state is still $\rho_{\text{ppm}}(0)=\rho_{S}\otimes\rho_{\text{vac}}$. The generator (24) exhibits the common feature of purified pseudomodes, namely, the annihilation and creation operators of any purified pseudomode always appear on the same side of the density matrix. Specifically, $d_{l+}$ and $d_{l+}^{\dagger}$ ($d_{l-}$ and $d_{l-}^{\dagger}$) only act on the left-hand (right-hand) side in Eq. (24). Since all purified pseudomodes evolve from their vacuum, this critical feature directly implies that the states of all purified pseudomodes can effectively be described in terms of pure states during the evolution.
@@
As a closing remark of this subsection, we note that by taking the limit $a\to+\infty$, the error related to the free parameters $a$ in the approximation (17) vanishes so that the only error source of the purified pseudomode model in Eq. (23) comes from the spectral decompositions in Eqs. (15) and (16). In addition, the equation of motion in Eq. (23) can be recast in the form of the HEOM, as shown explicitly in Appendix. B.
III.4 Purified pseudomode model for the nonlinear system-bath interaction (2)
@@
The purification procedure described in Appendix. III.3 is additive, in the sense that for the system-bath interaction $H_{\text{int}}=sX^{n}+sX^{m}$ the generator $\mathcal{L}_{\text{ppf}}$ of the corresponding purified pseudomode model is simply $\mathcal{L}_{\text{ppf}}=\mathcal{L}_{0}+\mathcal{L}_{\text{int},n}+\mathcal{L}_{\text{int},m}$. This fact means that, for the general nonlinear interaction (2), the corresponding generator can be directly written as $\mathcal{L}_{\text{ppf}}=\mathcal{L}_{0}+\sum_{n=1}^{\infty}\alpha_{n}\mathcal{L}_{\text{int},n}$. This concludes our construction of the purified pseudomode model.
@@
Figure 1: Spontaneous decay of a two-level atom placed inside a lossy cavity. Colorful solid curves are obtained from the corresponding purified pseudomode model (abbreviated as “PPM”) for four different combinations of the coefficients $(\alpha_{1},\alpha_{2},\alpha_{3})$, while black dashed lines are the results of the master equation in Eq. (25). Simulation parameters are $\omega_{S}=0.5$, $\nu=0.5$, $\gamma=0.1$, and $\lambda=0.3$.
[…]
@@ IV.1 Spontaneous decay of a two-level atom in a lossy cavity
where $s=\sigma_{x}$, $H_{S}=\omega_{S}\sigma_{z}/2$ is the bare atomic Hamiltonian with $\omega_{S}$ the level splitting, $b$ and $b^{\dagger}$ denote the annihilation and creation operators of the cavity, respectively. We consider a generalized light-matter interaction consisting of both the linear term and the second- and third-order corrections. In other words, the cavity coupling operator $Q_{\text{ac}}$ is written as $Q_{\text{ac}}=\alpha_{1}X_{\text{c}}+\alpha_{2}X_{\text{c}}^{2}+\alpha_{3}X_{\text{c}}^{3}$ with $X_{\text{c}}=\lambda(b+b^{\dagger})$ and $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{R}$. The lossy cavity can be viewed as the environment of the two-level atom and its free correlation can be solved analytically as $C(t)=\lambda^{2}\exp(-i\nu t-\gamma\lvert t\rvert)$. The model considered here can be engineered in superconducting quantum circuits [86, 87]. Since its bath correlation function is just a single exponential, no extra spectral decomposition is needed, which we take advantage of to demonstrate the exactness of the purification procedure described in Sec. III.3.
@@
In Fig. 1, we plot the dynamics of the atomic population and atom-cavity entanglement measured by the von-Neumann entropy for four different groups of the coefficients $(\alpha_{1},\alpha_{2},\alpha_{3})$ for a resonant cavity $\nu=\omega_{S}$. We compare the results obtained from the purified pseudomode model (solid lines) with those from the master equation in Eq. (25) (dashed black lines) and find full agreement in all cases. The presence of the small quadratic correction with $\alpha_{2}=0.4$ slightly increases the hybridization in the steady-state, while the the entanglement between the atom and the cavity decreases, as indicated by the red lines in Fig. 1. On the other hand, by tuning the sign of the cubic correction, stronger ($\alpha_{2}=0.15$, green lines) or weaker ($\alpha_{2}=-0.15$, purple lines) light-matter hybridization and entanglement can be engineered, possibly offering an extra control knob for quantum information processing and dissipative state engineering [88]. Note that, we have checked that the sign of the quadratic correction has no influence on the hybridization and entanglement (not shown).
@@
Despite the general validity of the purified pseudomode model constructed in Sec. III, its numerical simulations may encounter instability issues for large nonlinearity, which arises from the necessary truncation of the infinite-dimensional Hilbert spaces of purified pseudomodes. In fact, similar issues may also occur for pseudomode models involving unphysical model parameters and the HEOM, and several solutions to circumvent numerical instability have been proposed [89, 90] in the latter context. In our simulations, we have carefully chosen model parameters to avoid such instability. A more careful study of this practical issue will be left for future works.

### v2

@@ III.2 Explicit pseudomode model for the nonlinear system-bath interaction (2)
As in the analysis of Ref. [79], we assume a spectral decomposition of the positive contribution $C(t\geq 0)$ in the form of
@@
with $w_{l}\in\mathbb{C}$, $\nu_{l}\in\mathbb{R}$, $\gamma_{l}\in\mathbb{R}_{+}$, and the condition $\sum_{l=1}^{N}w_{l}\cong\sigma^{2}$ is satisfied for consistency. Such an exponential decomposition can be realized using several existing algorithms, e.g., the adaptive Antoulas-Anderson (AAA) algorithm and the Prony decomposition, see Ref. [82] and the references therein for a comprehensive review. Because of the time-reversal property $C(t)=C^{*}(-t)$, the decomposition in Eq. (15) also gives the decomposition of the negative contribution $C(t\leq 0)$, written as
@@
To match the pseudomode correlation ansatz (13), we combine the two decompositions in Eqs. (15) and (16) with the representation $\Theta(\pm t)\approx e^{\pm at-a\lvert t\rvert}$ (where $a\in[0,+\infty)$ is a sufficiently large free parameter) of the Heaviside step function to obtain the following approximation
@@
$C_{\text{pm},\gtrless}^{a}(t)=\sum_{l=1}^{N}(\text{Re}\,w_{l}\pm i\text{Im}\,w_{l})e^{-i(\nu_{l}\pm ia)t-(\gamma_{l}+a)\lvert t\rvert}$. (18)
The use of the symbol $\cong$ in Eq. (17) highlights that the approximating error is dominated by the one made using the approximation ansatz in Eq. (15) and it requires further justification, which we now provide. Comparing Eq. (18) with Eq. (15) and Eq. (16), we find that
[…]
@@
The collection of these parameters defines $2N+1$ zero-temperature pseudomodes whose correlation $C_{\text{pm}}^{a}(t)$ approximates $C(t)$ in the limit $a\to+\infty$. This completes the construction of an explicit pseudomode model for general nonlinear system-bath interactions (2). In the next section, we show that the modes constituting this model can be further purified to allow a more optimized numerical implementation of the corresponding algorithm.
III.3 Purified pseudomode model for the nonlinear system-bath interaction (2)
In our previous work [79] where linear system-bath interactions were considered, we have shown that under the analytical continuation in Eq. (21) for the pseudomode energy and decay parameters, the state of the pseudomodes becomes effectively pure in the limit $a\to+\infty$, leading to an optimization of the pseudomode Hilbert space which can be exploited in numerical simulations. The modes associated to these analytically-continued parameters are denoted as “purified pseudomodes”. Here we extend the results in Ref. [79] to incorporate nonlinear interactions.
@@ III.3 Purified pseudomode model for the nonlinear system-bath interaction (2)
In Appendix. A, we show that the purification technique developed in Ref. [79] can be generalized to the case $Q=X^{n}$. Specifically, we show that, in the limit $a\to+\infty$, the equation of motion in Eq. (10) becomes the following Lindblad-like master equation
$i\frac{d\rho_{\text{ppm}}(t)}{dt}=\mathcal{L}_{S}\rho_{\text{ppm}}+\mathcal{L}_{0}\rho_{\text{ppm}}+\mathcal{L}_{\text{int},n}\rho_{\text{ppm}},$ (23)
where $\mathcal{L}_{S}[\boldsymbol{\cdot}]=[H_{S},\boldsymbol{\cdot}]$ and the generators $\mathcal{L}_{0}$ and $\mathcal{L}_{\text{int},n}$ are defined as
@@
$\begin{array}[]{lll}&\displaystyle\mathcal{L}_{0}[\boldsymbol{\cdot}]=\sum_{l=1}^{N}\Big[(\nu_{l}-i\gamma_{l})d_{l+}^{\dagger}d_{l+}[\boldsymbol{\cdot}]-(\nu_{l}+i\gamma_{l})[\boldsymbol{\cdot}]d_{l-}^{\dagger}d_{l-}\Big],\\ &\displaystyle\mathcal{L}_{\text{int},n}[\boldsymbol{\cdot}]=X_{+}^{n}s[\boldsymbol{\cdot}]-[\boldsymbol{\cdot}]sX_{-}^{n}+s[\boldsymbol{\cdot}]X_{-,d^{\dagger}}^{n}-X_{+,d}^{n}[\boldsymbol{\cdot}]s\\ &\displaystyle~~~~~~~~~~~+\sum_{r=1}^{n-1}\binom{n}{r}\bigg[X_{+}^{n-r}s[\boldsymbol{\cdot}]X_{-,d^{\dagger}}^{r}-X_{+,d}^{r}[\boldsymbol{\cdot}]sX_{-}^{n-r}\bigg].\end{array}$ (24)
@@
Here, $\rho_{\text{ppm}}(t)$ denotes the density matrix in the limit $a\to+\infty$, and we have defined the operators $X_{\pm}=X_{\pm,d}+X_{\pm,d^{\dagger}}$ with $X_{\pm,d}=\sum_{l=1}^{N}\lambda_{l\pm}d_{l\pm}$ and $X_{\pm,d^{\dagger}}=\sum_{l=1}^{N}\lambda_{l\pm}d_{l\pm}^{\dagger}$ their components to simplify the notation. The initial state is still $\rho_{\text{ppm}}(0)=\rho_{S}\otimes\rho_{\text{vac}}$. The generator (24) explicitly exhibits the main feature of purified pseudomodes, namely that the annihilation and creation operators of any purified pseudomode always appear on the same side of the density matrix. Specifically, $d_{l+}$ and $d_{l+}^{\dagger}$ ($d_{l-}$ and $d_{l-}^{\dagger}$) only act on the left-hand (right-hand) side in Eq. (24). Since all purified pseudomodes evolve from their vacuum, this critical feature directly implies that the states of all purified pseudomodes can effectively be described in terms of pure states during the evolution.
As an important remark, we note that by taking the limit $a\to+\infty$ in Eq. (18), the free correlations of the purified pseudomodes $d_{l\pm}$ can be formally identified as
$C_{\text{ppm}}^{l\pm}(t)=(\text{Re}\,w_{l}\pm i\text{Im}\,w_{l})e^{-i\nu_{l}t-\gamma_{l}t}\Theta(\pm t).$ (25)
@@
In addition, the error related to the free parameters $a$ in the approximation (17) vanishes in this limit so that the only error source of the purified pseudomode model in Eq. (23) comes from the spectral decompositions in Eqs. (15) and (16). As a side remark, we also note that the equation of motion in Eq. (23) can be recast in the form of the HEOM, as shown explicitly in Appendix. B.
@@
We have now all the ingredients in order to obtain the purified model for the full interaction in Eq. (2). In fact, as it can be explicitly seen in Appendix. A, the procedure to derive the model in Eq. (23) is additive, in the sense that for a system-bath interaction $H_{\text{int}}=sX^{n}+sX^{m}$ the generator $\mathcal{L}_{\text{ppm}}$ of the corresponding purified pseudomode model is simply $\mathcal{L}_{\text{ppm}}=\mathcal{L}_{\text{S}}+\mathcal{L}_{0}+\mathcal{L}_{\text{int},n}+\mathcal{L}_{\text{int},m}$. This additivity arises from the fact that the coupling operator $Q_{\text{pm}}$ in Eq. (14) is expressed as the sum of all nonlinear contributions $X_{\text{pm}}^{n}$ with weights $\alpha_{n}$, and the fact that the limit operation is additive. As a result, for the general nonlinear interaction (2), the corresponding generator can be directly written as
$\mathcal{L}_{\text{ppm}}=\mathcal{L}_{S}+\mathcal{L}_{0}+\sum_{n=1}^{\infty}\alpha_{n}\mathcal{L}_{\text{int},n}\;,$ (26)
thereby concluding our construction of the purified pseudomode model.
@@
Figure 1: Spontaneous decay of a two-level atom placed inside a lossy cavity. Panel (a) and (b) show the evolution of the excited state population $\langle\sigma_{+}\sigma_{-}\rangle$ and the von Neumann entropy $S_{\text{vN}}$, respectively. Colored solid curves are obtained using the corresponding purified pseudomode model (abbreviated as “PPM”) for four different combinations of the coefficients $(\alpha_{1},\alpha_{2},\alpha_{3})$ in the coupling operator $Q_{\text{ac}}=\alpha_{1}X_{\text{c}}+\alpha_{2}X_{\text{c}}^{2}+\alpha_{3}X_{\text{c}}^{3}$ with $X_{\text{c}}=\lambda(b+b^{\dagger})$. Black dashed lines are the solution of the master equation in Eq. (27). Simulation parameters are $\omega_{S}=0.5$, $\nu=0.5$, $\gamma=0.1$, and $\lambda=0.3$.
Figure 2: Excited state population $\langle\sigma_{+}\sigma_{-}\rangle$ in panel (a) and von Neumann entropy $S_{\text{vN}}$ in panel (b) for the steady-state of the two-level atom placed in a lossy cavity. Here $\alpha_{1}=1$ is fixed in the coupling operator $Q_{\text{ac}}=\alpha_{1}X_{\text{c}}+\alpha_{2}X_{\text{c}}^{2}+\alpha_{3}X_{\text{c}}^{3}$. Other simulation parameters are the same as those in Fig. 1.
Figure 3: (a) Power spectrum $S(\omega)$ (blue solid) of the phonon environment and its approximation (red dashed) generated with the AAA algorithm. The positive (negative) frequency domain of $S(\omega)$ corresponds to emission (absorption) of phonons, respectively. The upper inset shows the fitting error $\epsilon_{S}=\int d\omega\lvert S(\omega)-S_{\text{AAA}}(\omega)\rvert$ for different NN, while the lower inset shows the fitting error $\epsilon_{C}=\lvert C(t)-C_{\text{AAA}}(t)\rvert$ as a function of time $t$. (b) Resonance fluorescence spectra in the absence (blue line) and presence (orange line for the pure linear coupling and green line for the linear+quadratic coupling) of the phonon environment. Other simulation parameters are $\kappa=$0.1\text{\,}\mathrm{meV}$\$ and $\Omega=$1\text{\,}\mathrm{meV}$\$.
@@ IV.1 Spontaneous decay of a two-level atom in a lossy cavity
[…]
The model considered here, i.e., the nonlinear open Rabi model, can be engineered in superconducting quantum circuits [86, 87]. Nonlinear light-matter interactions, such as the two-photon coupling considered in Refs. [86, 87, 88], give rise to a number of interesting phenomena, for example, spectral collapse [89, 90], high-order optical nonlinear processes [91, 92], and novel collective photon emission [93, 94]. They could also find applications in quantum technologies, such as quantum sensing [95], qubit-noise spectroscopy [96] and nonclassical state preparation [97].
@@
In Fig. 1, we plot the evolution of the excited state population $\langle\sigma_{+}\sigma_{-}\rangle$ and atom-cavity entanglement measured by the von-Neumann entropy $S_{\text{vN}}=-\text{Tr}[\rho_{S}\ln\rho_{S}]$ for four different groups of the coefficients $(\alpha_{1},\alpha_{2},\alpha_{3})$ for a resonant cavity $\nu=\omega_{S}$. Note that, the von-Neumann entropy is a valid entanglement measure only for zero-temperature environments [98], which is the case considered here. We compare the results obtained from the purified pseudomode model (solid lines) with those from the master equation in Eq. (27) (dashed black lines) and find full agreement in all cases. The presence of the small quadratic correction with $\alpha_{2}=0.4$ slightly increases the hybridization in the steady-state, while the the entanglement between the atom and the cavity decreases, as indicated by the red lines in Fig. 1. On the other hand, by tuning the sign of the cubic correction, stronger ($\alpha_{3}=0.15$, green lines) or weaker ($\alpha_{3}=-0.15$, purple lines) light-matter hybridization and entanglement can be engineered, possibly offering an extra control knob for quantum information processing and dissipative state engineering [99]. In Fig. 2 we further plot $\langle\sigma_{+}\sigma_{-}\rangle$ and $S_{\text{vN}}$ as functions of $\alpha_{2}$ and $\alpha_{3}$ with fixed $\alpha_{1}=1$ for the steady-state, which clearly reveals the opposite dependence of the excited-state population and entropy on the sign of $\alpha_{3}$, in stark contrast to their dependence on $\alpha_{2}$. Note that, we have checked that the sign of the quadratic correction has no influence on the hybridization and entanglement (not shown).
@@
Despite the general validity of the purified pseudomode model constructed in Sec. III, its numerical simulations may encounter instability issues for large nonlinearity, which arises from the necessary truncation of the infinite-dimensional Hilbert spaces of purified pseudomodes. Specifically, numerical simulations of the master equation in Eq. (23) require truncating the infinite-dimensional pseudomode Hilbert space to a finite dimension $M$. As a consequence, the corresponding truncated generator $\mathcal{L}_{\text{ppm}}^{(M)}$ may give rise to eigenvalues $\xi_{j}^{(M)}$ with large positive imaginary part, $\text{Im}\,\xi_{j}^{(M)}>0$, even if the fitting error $\lvert C(t)-C_{\text{pm}}^{\infty}(t)\rvert$ is sufficiently small. Even worse, in certain circumstances $\text{Im}\,\xi_{j}^{(M)}$ decrease very slowly as $M$ increases. This leads to divergence in the dynamics, preventing the numerical evaluation of the steady-state using the truncated generator $\mathcal{L}_{\text{ppm}}^{(M)}$. In fact, similar issues may also occur for other pseudomode models involving unphysical model parameters and the HEOM, and several solutions to circumvent numerical instability have been proposed [100, 101] in the latter context. In our simulations, we have carefully chosen model parameters to avoid such instability. A more careful study of this practical issue will be left for future works.